OpenStudy (anonymous):

Find the slope of the line that passes through the points (2,-4), and (4,8).

OpenStudy (cornzyblack):

y2 = 8, y1 = -4 x2 = 4, x1 = 2 using m = (y2 - y1)/(x2 - x1) = (8 - (-4))/(4 - 2) = 12/2 = 6